OpenStudy (anonymous):

Please Help:(( Evaluate the following expressions. Given F(x) = x2 + 2, G(x) = 3x + 1, H(x) = x F(2) =

OpenStudy (ksaimouli):

(2)^2+2 =? and plz vote http://www.signazon.com/contest/grad2014/abgrqq88/

OpenStudy (anonymous):

so its 6?

OpenStudy (zzr0ck3r):

\(f(x) = x^2+2 \implies \ f(2) = 2^2+2 = 6 \)

OpenStudy (zzr0ck3r):

correct

OpenStudy (ksaimouli):

yup